# Models of Set Theory: Extensions and Dead-ends

### v5

@@ 3. TALLER CONSERVATIVE ELEMENTARY EXTENSIONS
Thus $\left(\mathcal{M},\Phi\cap M\right)$ satisfies $w\neq\varnothing\wedge\forall t\in w\exists t^{\prime}\in w\wedge\left((t^{\prime}\in t)\in\Phi\right).$ $\mathcal{N}$ can evaluate each term $t(c)$ in $w$ as an ordinal $\delta<\beta$, where $\mathcal{N}_{\beta}\models t(c_{\lambda})=\delta$ (here $t(c_{\lambda})$ is the term obtained by replacing all occurrences of the constant $c$ with $c_{\lambda}).$ So we can consider $s\in N$, where:
$\mathcal{N}\models s=\left\{\delta:\exists t(c)\in w:(\mathrm{V}_{\beta},\in)\models t(c_{\lambda}):t(c)\in w\right\}.$
Within $\mathcal{N}$, $s$ is a nonempty of ordinals that has no least element, which of course is a contradiction. This shows that Case II is impossible, thus concluding the proof. $\square$
4.5. Remark. Arguing in $\mathrm{ZFC}$, suppose $\kappa$ is the first strongly inaccessible cardinal, $\lambda$ is the second strongly inaccessible cardinal, $\mathcal{M}=(\mathrm{V}_{\kappa},\in)$ and $\mathcal{N}=(\mathrm{V}_{\lambda},\in)$. Then $\mathcal{N}$ is a proper faithful rank extension of $\mathcal{M}$ in which the first clause of the dichotomy of the conclusion of part (a) of Theorem 4.1 (and 4.4) holds , but not the second one. On the other hand, if $\kappa$ is a weakly compact cardinal, then it is well-known that $(\mathrm{V}_{\kappa},\in)$ has a faithful topless elementary end extension $\mathcal{N}$; in this scenario the second clause of the dichotomy of the conclusion of part (a) of Theorem 4.1 (and 4.4) holds, but not the first one.
4.6. Remark. If the assumptions of Theorem 4.4 are strengthened by adding the assumption that $\mathcal{M}\prec\mathcal{N}$, then as in the proof of Theorem 3.3 of [En-3] we can use ‘Kaufmann’s trick’ to conclude that there is some $\mathcal{K}$ such that $\mathcal{M}\prec\mathcal{K}\preceq\mathcal{N}$ and $\mathcal{K}$ is a topped rank extension of $\mathcal{M}$. This yields a strengthening of Theorem 3.3 of [En-3] by eliminating the assumption that the axiom of choice is true in $\mathcal{M}$ and $\mathcal{N}$.
4.7. Remark. An inspection of the proof of Theorem 4.1 makes it clear that the assumption that $\mathcal{M}\models\mathrm{ZFC}$ can be reduced to $\mathcal{M}\models\mathrm{ZF+DC}$, and the assumption that $\mathcal{N}\models\mathrm{ZFC}$ can be reduced to $\mathcal{N}\models\mathrm{KPR}$ + “every set can be well-ordered”. Here $\mathrm{KPR}$ is as in part (d) of Definition 2.1. Similarly, an inspection of the proof of Theorem 4.4 shows that the assumption that $\mathcal{N}\models\mathrm{ZF}$ can be reduced to $\mathcal{N}\models\mathrm{KPR}$.
4.8. Question. Can Theorem 4.1 and Corollary 4.2 be strengthened by only assuming that $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathrm{ZF}$?
@@
In this section we will establish Theorem C of the abstract, the proof uses many ingredients, including the following one that refines a result obtained independently by Kaufmann and the author who demonstrated Theorem 5.1 for models of $\mathrm{ZF}$ in which the Axiom of Choice holds (see “Added in Proof” of [Kau] and Remark 1.6 of [En-2]).
5.1. Theorem. No model of $\mathrm{ZF}$ has a conservative proper end extension satisfying $\mathrm{ZF}$.
Proof. This follows from putting part (b) of Theorem 4.4 together with Tarski’s Undefinability of Truth Theorem (as in Theorem 2.8). $\square$.□ .
@@
(a) $X\subseteq M$ is piecewise-coded in $\mathcal{M}$ if $\forall a\in M\ \exists b\in M\ \mathrm{Ext}_{\mathcal{M}}(b)=X\cap\mathrm{Ext}_{\mathcal{M}}(a).$
(b) $\mathcal{M}$ is rather classless iff every piecewise-coded subset of $M$ is $\mathcal{M}$-definable.
(c) $\mathcal{M}$ is $\aleph_{1}$-like if $\left|M\right|=\aleph_{1}$ but $\left|\mathrm{Ext}_{\mathcal{M}}(a)\right|\leq\aleph_{0}$ for each $a\in M.$
5.3. Theorem. (Keisler-Kunen [Ke-1], Shelah [Sh]). Every countable model of $\mathrm{ZF}$ has an elementary end extension to an $\aleph_{1}$-like rather classless model.
5.4. Proposition. No rather classless model of $\mathrm{ZF}$ has a proper rank extension to a model of $\mathrm{ZF}$.
Proof. This follows from putting Theorem 4.4 together with the observation that a rank extension of a model of $\mathrm{ZF}$ is a conservative extension, and therefore a faithful extension. $\square$
5.5. Remark. As pointed out in Remark 1.6 of [En-3] it is possible for a rather classless model to have a proper end extension satisfying $\mathrm{ZF}$, since $\aleph_{1}$-like rather classless models exist by Theorem 5.3, and one can use the Boolean-valued approach to forcing to construct set generic extensions of such models.
@@
(1) $(\mathbb{T},\ \leq_{\mathbb{T}})$ is a tree, i.e., a partial order such that any two elements below a given element are comparable.
(2) $(\mathbb{L},\ \leq_{\mathbb{L}})$ is a linear order with no last element.
(3) $\rho$ is an order preserving map from $(\mathbb{T},\ \leq_{\mathbb{T}})$ onto $(\mathbb{L},\ \leq_{\mathbb{L}})$ with the property that for each $t\in\mathbb{T},$ $\rho$ maps the set of predecessors of $t$ onto the initial segment of $(\mathbb{L},\ \leq_{\mathbb{L}})$ consisting of elements of $L$ less than $\rho(t).$

### v6

@@ 3. $(iii)$
Thus $\left(\mathcal{M},\Phi\right)\models$ $w\neq\varnothing\wedge\exists t^{\prime}\in\Gamma\ (t^{\prime}\in t)\in\Phi.$ Observe that since $\mathcal{N}$ has access to the Tarskian satisfaction class for $\mathcal{N}_{\beta}$, $\mathcal{N}$ can evaluate each term $t(c)$ in $w$ as an ordinal $\delta<\beta$, where $\mathcal{N}_{\beta}\models t(\dot{\lambda})=\delta$, where $t(\dot{\lambda})$ is the term obtained by replacing all occurrences of the constant $c$ with $\dot{\lambda}$ in $t(c).$ So we can consider $s\in N$, where:
$\mathcal{N}\models s=\left\{\delta:\exists t(c)\in w\ (\mathrm{V}_{\beta},\in)\models\delta=t(\dot{\lambda})\right\}.$
Within $\mathcal{N}$, $s$ is a nonempty set of ordinals that has no least element, which of course is a contradiction. This shows that Case II is impossible, thus concluding the proof. $\square$
4.5. Remark. Arguing in $\mathsf{ZFC}$, suppose $\kappa$ is the first strongly inaccessible cardinal, $\lambda$ is the second strongly inaccessible cardinal, $\mathcal{M}=(\mathrm{V}_{\kappa},\in)$ and $\mathcal{N}=(\mathrm{V}_{\lambda},\in)$. Then $\mathcal{N}$ is a proper faithful rank extension of $\mathcal{M}$ in which the first clause of the dichotomy of the conclusion of part (a) of Theorem 4.1 (and 4.4) holds , but not the second one. On the other hand, if $\kappa$ is a weakly compact cardinal, then as noted by Kaufmann [Kau, Proposition 2.3], $(\mathrm{V}_{\kappa},\in)$ has a faithful topless elementary end extension $\mathcal{N}$; in this scenario the second clause of the dichotomy of the conclusion of part (a) of Theorem 4.1 (and 4.4) holds, but not the first one.
4.6. Remark. If the assumptions of Theorem 4.4 are strengthened by adding the assumption that $\mathcal{M}\prec\mathcal{N}$, then as in the proof of Theorem 3.3 of [En-3] we can use ‘Kaufmann’s trick’ to conclude that there is some $\mathcal{K}$ such that $\mathcal{M}\prec\mathcal{K}\preceq\mathcal{N}$ and $\mathcal{K}$ is a topped rank extension of $\mathcal{M}$. This yields a strengthening of Theorem 3.3 of [En-3] by eliminating the assumption that the axiom of choice holds in $\mathcal{M}$ and $\mathcal{N}$.
4.7. Remark. An inspection of the proof of Theorem 4.1 makes it clear that the assumption that $\mathcal{M}\models\mathsf{ZFC}$ can be reduced to $\mathcal{M}\models\mathsf{ZF}+\mathsf{DC}$, and the assumption that $\mathcal{N}\models\mathsf{ZFC}$ can be reduced to $\mathcal{N}\models\mathsf{KPR}$ + “every set can be well-ordered”. Here $\mathsf{KPR}$ is as in part (d) of Definition 2.1. Similarly, an inspection of the proof of Theorem 4.4 shows that the assumption that $\mathcal{N}\models\mathsf{ZF}$ can be reduced to $\mathcal{N}\models\mathsf{KPR}$.
4.8. Question. Can Theorem 4.1 and Corollary 4.2 be strengthened by assuming that $\mathcal{M}$ and $\mathcal{N}$ are models of $\mathsf{ZF}$?
@@
In this section we will establish Theorem C of the abstract, the proof uses many ingredients, including the following one that refines a result obtained independently by Kaufmann and the author who demonstrated Theorem 5.1 for models of $\mathsf{ZF}$ in which the Axiom of Choice holds (see “Added in Proof” of [Kau] and Remark 1.6 of [En-2]).
5.1. Theorem. No model of $\mathsf{ZF}$ has a conservative proper end extension satisfying $\mathsf{ZF}$.
Proof. This follows from putting part (b) of Theorem 4.4 together with Tarski’s Undefinability of Truth Theorem (as in Theorem 2.8). $\square.$
@@
  1. (a)
    $X\subseteq M$ is a class [^29] Classes of $\mathcal{M}$ are sometimes referred to as piecewise coded subsets of $\mathcal{M}$. in $\mathcal{M}$ if $\forall a\in M\ \exists b\in M\ X\cap\mathrm{Ext}_{\mathcal{M}}(a)=\mathrm{Ext}_{\mathcal{M}}(b).$
  2. (b)
    $\mathcal{M}$ is rather classless if every class of $M$ is $\mathcal{M}$-definable.
  3. (c)
    $\mathcal{M}$ is $\aleph_{1}$-like if $\left|M\right|=\aleph_{1}$ but $\left|\mathrm{Ext}_{\mathcal{M}}(a)\right|\leq\aleph_{0}$ for each $a\in M.$
5.3. Theorem. (Keisler-Kunen [Ke-1], Shelah [Sh]). Every countable model of $\mathsf{ZF}$ has an elementary end extension to an $\aleph_{1}$-like rather classless model.
5.4. Proposition. No rather classless model of $\mathsf{ZF}$ has a proper rank extension to a model of $\mathsf{ZF}$.
Proof. This follows from putting Theorem 4.4 together with the observation that a rank extension of a rather classless model of $\mathsf{ZF}$ is a conservative extension, and therefore a faithful extension. $\square$
5.5. Remark. As pointed out in Remark 1.6 of [En-3] it is possible for a rather classless model to have a proper end extension satisfying $\mathsf{ZF}$, since $\aleph_{1}$-like rather classless models exist by Theorem 5.3, and one can use the Boolean-valued approach to forcing to construct set generic extensions of such models.
@@
  1. (1)
    $(\mathbb{T},\ \leq_{\mathbb{T}})$ is a tree, i.e., a partial order such that any two elements below a given element are comparable.
  2. (2)
    $(\mathbb{L},\ \leq_{\mathbb{L}})$ is a linear order with no last element.
  3. (3)
    $\rho$ is an order preserving map from $(\mathbb{T},\ \leq_{\mathbb{T}})$ onto $(\mathbb{L},\ \leq_{\mathbb{L}})$ with the property that for each $t\in\mathbb{T},$ $\rho$ maps the set of predecessors of $t$ onto the initial segment of $(\mathbb{L},\ \leq_{\mathbb{L}})$ consisting of elements of $L$ less than $\rho(t).$
@@
5.8. Definition. Given a structure $\mathcal{M}$ in a language $\mathcal{L}$, we say that a ranked tree $\tau$ is $\mathcal{M}$-definable if $\tau=\mathbf{t}^{\mathcal{M}}$, where $\mathbf{t}$ is an appropriate sequence of $\mathcal{L}$-formulae whose components define the corresponding components of $\tau$ in $\mathcal{M}.$ $\mathcal{M}$ is rather branchless if for each $\mathcal{M}$-definable ranked tree $\tau$, all branches of $\tau$ (if any) are $\mathcal{M}$-definable.
  1. (a)
    (Keisler-Kunen [Ke-1], essentially). It is a theorem of $\mathsf{ZFC}+\Diamond_{\omega_{1}}$ that $\mathcal{M}$ can be elementarily extended to a rather branchless model.
  2. (b)
    (Shelah [Sh]). It is a theorem of $\mathsf{ZFC}$ that $\mathcal{M}$ can be elementarily extended to a rather branchless model.
  1. (a)
    $(\mathbb{P},\leq_{\mathbb{P}})$ is directed if any given pair of elements of $\mathbb{P}$ has an upper bound.
  2. (b)
    A subset $F$ of $\mathbb{P}$ is a filter over $(\mathbb{P},\leq_{\mathbb{P}})$ if the sub-poset $(F,\leq_{\mathbb{P}})$ is directed.
  3. (c)
    A filter over $(\mathbb{P},\leq_{\mathbb{P}})$ is maximal if it cannot be properly extended to a filter over $(\mathbb{P},\leq_{\mathbb{P}})$. [^30] Thus a maximal filter can be described as a maximally compatible subset of $(\mathbb{P},\leq_{\mathbb{P}}).$
  4. (d)
    A subset $C$ of $\mathbb{P}$ is cofinal in $(\mathbb{P},\leq_{\mathbb{P}})$ if $\forall x\in\mathbb{P}\ \exists y\in C\ x\leq_{\mathbb{P}}y.$
5.11. Definition. Suppose $s$ is an infinite set.
  1. (a)
    $s]^{<\omega}$ is the directed poset of finite subsets of $s,$ ordered by containment. Note that $s]^{<\omega}$ is a directed set with no maximum element.
  2. (b)
    $\mathrm{Fin}(s,2)$ is the poset of finite functions from $s$ into $\{0,1\}$, ordered by containment (where a function is viewed as a set of ordered pairs).
5.12. Example. Given an infinite set $a$, and $s\subseteq a,$ let $\chi_{s}:s\rightarrow 2$ be the characteristic function of $s,$ i.e., $\chi_{s}(x)=1$ iff $x\in s.$ Clearly $\chi_{s}]^{<\omega}$ is a maximal filter of $\mathrm{Fin}(a,2).$ More generally, if $\mathcal{N}\models\mathsf{ZF}$, and $\mathcal{N}\modelssas\subseteq$ a and $a$ is infinite”, consider $(\mathbb{P},\leq_{\mathbb{P}})$, where:
$\mathbb{P}:=\mathrm{Ext}_{\mathcal{N}}(\mathrm{Fin}^{\mathcal{N}}(a,2))$,
and $\leq_{\mathbb{P}}$ is set-inclusion (among members of $\mathbb{P}$) in the sense of $\mathcal{N}$. Then $F_{s}$ is a maximal filter over $(\mathbb{P},\leq_{\mathbb{P}})$, where:
  1. (a)
    Every $\mathcal{M}$-definable directed poset $(\mathbb{D},\leq_{\mathbb{D}})$ with no maximum element has a cofinal chain of length $\omega_{1}.$
  2. (b)
    Given any $\mathcal{M}$-definable poset $\mathbb{P}$, and any maximal filter $F$ of $\mathbb{P}$, if $F$ has a cofinal chain of length $\omega_{1}$, then $F$ is coded in $\mathcal{M}$.
5.14. Remark. If $\tau=(\mathbb{T},\ \leq_{\mathbb{T}},\ \mathbb{L},\ \leq_{\mathbb{L}},\ \rho)$ is a ranked tree, then each branch of $\tau$ is a maximal filter over $(\mathbb{T},\ \leq_{\mathbb{T}}).$ This makes it clear that every Rubin model is rather branchless. Also, a rather branchless model of $\mathsf{ZF}$ is rather classless. To see this consider the ranked tree $\tau_{\mathrm{class}}$ defined within a model $\mathcal{M}$ of $\mathsf{ZF}$ as follows: The nodes of $\tau_{\mathrm{class}}$ are ordered pairs $(s,\alpha)$, where $s\subseteq\mathrm{V}_{\alpha}$, the rank of $(s,\alpha)$ is $\alpha$ and $(s,\alpha)<(t,\beta)$ if $\alpha\in\beta$ and $s=t\cap\mathrm{V}_{\alpha}$. It is easy to see that $\mathcal{M}$ is rather classless iff every branch of $\tau_{\mathrm{class}}^{\mathcal{M}}$ is $\mathcal{M}$-definable. [^31] The ranked tree $\tau_{\mathrm{class}}$ was first introduced by Keisler [Ke-1, Example 2.1], who noted that $\mathcal{M}$ is rather classless iff every branch of $\tau_{\mathrm{class}}^{\mathcal{M}}$ is $\mathcal{M}$-definable, which is the key property that we need here. As noted by the referee, this key property is also satisfied by the subtree $\sigma_{\mathrm{class}}$ of $\tau_{\mathrm{class}}$ whose tree-nodes are of the form $\left(\alpha,s\right)$ where the ordinal rank of $s$ is $\alpha.$ Note that the branches of $\tau_{\mathrm{class}}^{\mathcal{M}}$ correspond to all classes of $\mathcal{M}$ (including improper ones, i.e., those that form a set), whereas the branches of $\sigma_{\mathrm{class}}^{\mathcal{M}}$ correspond to all proper classes of $\mathcal{M}$. Hence we have the following chain of implications:
Rubin $\Rightarrow$ rather branchless $\Rightarrow$ rather classless.
5.15. Theorem. (Rubin [^32] The attribution to Rubin is informed by fact that the credit for [RS, Lemma 2.3] is explicitly given to Rubin (at the end of the introduction of the paper). [RS, Corollary 2.4]). It is a theorem of $\mathsf{ZFC}+\Diamond_{\omega_{1}}$ that if $\mathcal{M}$ is a countable structure in a countable language, then $\mathcal{M}$ has an elementary extension of cardinality $\aleph_{1}$ that is a Rubin model. [^33] The proof of Theorem 5.15 in [RS, Lemma 2.3] has a small gap (it is assumed that $T_{0}\vdash\exists\vec{x}\psi(\vec{x})$ instead of assuming the consistency of $T_{0}+\exists\vec{x}\psi(\vec{x})$). The improved proof was presented in [En-1, Theorem 2.1.3] and is reproduced here in the Appendix.
5.16. Definition. Suppose $\mathcal{M}$ is a model of $\mathsf{ZF}$. $\mathcal{M}$ is weakly Rubin if (a) and (b) below hold:
  1. (a)
    $\mathcal{M}$ is rather classless (this is the asymptotic case of $(ii)$ below, if one could use $a=\mathrm{V}$).
  2. (b)
    For every element $a$ of $\mathcal{M}$ that is infinite in the sense of $\mathcal{M}$ the following two statements hold:
    1. $(1)$
      $\left([a]^{<\omega}\right)^{\mathcal{M}}$ has a cofinal chain of length $\omega_{1}.$
    2. $(2)$
      If $F$ is a maximal filter of $\mathrm{Fin}^{\mathcal{M}}(a,2)$ and $F$ has a cofinal chain of length $\omega_{1}$, then $F$ is coded in $\mathcal{M}$.
As demonstrated in the Appendix, Schmerl’s strategy of $\Diamond_{\omega_{1}}$-elimination in [Schm-2] (which is based on an absoluteness argument first presented by Shelah [Sh]) can be employed to build weakly Rubin models within $\mathsf{ZFC}$. Thus we have:
5.17. Theorem (Rubin-Shelah-Schmerl) It is a theorem of $\mathsf{ZFC}$ that every countable model of $\mathsf{ZF}$ has an elementary extension to a weakly Rubin model of cardinality $\aleph_{1}$.
5.18. Theorem. Every countable model $\mathcal{M}_{0}\models\mathsf{ZF}$ has an elementary extension $\mathcal{M}$ of power $\aleph_{1}$ that has no proper end extension to a model $\mathcal{N}\models\mathsf{ZF}$. Thus every consistent extension of $\mathsf{ZF}$ has a model of power $\aleph_{1}$ that has no proper end extension to a model of $\mathsf{ZF}$.
Proof [^34] The proof of Fact $\left(\nabla\right)$ presented here is a reformulation of a proof that was suggested by the referee. Our original proof admittedly used too much machinery.. We begin with a basic fact that will be called upon in the proof.
Fact $\left(\nabla\right).$ Suppose $\mathcal{M}\models\mathsf{ZF}$ and $\mathcal{N}\models\mathsf{ZF}$ with $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$, and $a\in M$. If $\in N, ss$ is finite as viewed in $\mathcal{N}$, and $\mathcal{N}\models s\subseteq a$, then $s\in M.$ Thus for all $a\in M$, we have:
Proof. The assumptions on $\mathcal{M}$ and $\mathcal{N}$ readily imply:
(1) $\omega^{\mathcal{M}}=\omega^{\mathcal{N}}$, and
(2) $\mathcal{M}\preceq_{\Delta_{1}}\mathcal{N}$.
Let $x=\mathrm{Fn}(y,a)$ be the usual formula expressing “$x$ is the set of all functions $f:y\rightarrow a$. It is easy to see that:
(3) The predicate $\left(x=\mathrm{Fn}(y,a)\wedge y\in\omega\right)$ is $\Delta_{1}$ within $\mathsf{ZF.}$
Observe that $x=\mathrm{Fn}(y,a)$ is clearly expressible by a $\Pi_{1}$-formula within $\mathsf{ZF}$ (with no restriction on $y$). With the added condition that $y\in\omega$, we can take advantage of recursion to express $x=\mathrm{Fn}(y,a)$ by the following $\Sigma_{1}$-formula:
$\exists\left\langle s_{0},\cdot\cdot\cdot,s_{y}\right\rangle$ $\left[s_{y}=x\wedge s_{0}=\{\varnothing\}\wedge\forall z<y\ (s_{z+1}=\left\{f\cup\{(z,v)\}:f\in s_{z}\wedge v\in a\right\})\right].$
Now assume $s\in N$, $\mathcal{N}\models s\subseteq a$, and for some $k\in N,$ $\mathcal{N\models}$ $k\in\omega\wedge\left|s\right|=k$. Thanks to (1) and the assumption $\mathcal{M}\subseteq_{\mathrm{end}}\mathcal{N}$, $k\in M.$ Therefore by (2) and (3), if $b\in M$ such that $\mathcal{M}\models b=\mathrm{Fn}(k,a)$, then $\mathcal{N}\models b=\mathrm{Fn}(k,a)$. Thus if $f\in N$ such that $f:k\rightarrow a$ is an injective function such that $\mathrm{range}(f)=s$, then $f\in b$, and since $\mathcal{N}$ is an end extension of $\mathcal{M}$, $f\in M.$ Hence $\mathrm{range}(f)=s\in M.$ $\square$ Fact $\left(\nabla\right)$)
Given a countable model $\mathcal{M}_{0}$ of $\mathsf{ZFC}$, by Theorem 5.17 there is a weakly Rubin model $\mathcal{M}$ that elementarily extends $\mathcal{M}_{0}$. By Theorem 5.1, to prove Theorem 5.18 it is sufficient to verify that every end extension $\mathcal{N}$ of $\mathcal{M}$ that satisfies $\mathsf{ZF}$ is a conservative extension. Towards this goal, suppose $\mathcal{M}\subsetneq_{\mathrm{end}}\mathcal{N}\models\mathsf{ZF}$. In light of Remark 2.6(e), it suffices to show that $\mathcal{M}\subsetneq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$.
To show $\mathcal{M}\subsetneq_{\mathrm{end}}^{\mathcal{P}}\mathcal{N}$, suppose that for $a\in M$ and $s\in N$, $\mathcal{N}\models s\subseteq a.$ We will show that $s\in M.$ By Fact $(\nabla)$ we may assume that $a$ is infinite as viewed from $\mathcal{M}$. Note that this implies that $\mathcal{M}$ views $a]^{<\omega}$ as a directed set with no maximum element. Fact $(\nabla)$ assures us that: